The measure of each of the congruent angles of an isosceles triangle is 9 less than 4 times the vertex angle. Find the measure of each angle of the triangle.
Answer by lwsshak3(11628)

The measure of each of the congruent angles of an isosceles triangle is 9 less than 4 times the vertex angle. Find the measure of each angle of the triangle.
**
let x=vertex
4x-9=congruent angle
..
vertex+2*congruent angles=180
x+2(4x-9)=180
x+8x-18=180
9x=198
x=198/9=22º
4x-9=88-9=79º
ans:
vertex=22º
each congruent angle=79º